40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4962     Entry time: Tue Jul 12 11:52:54 2011
Author: Jamie, Suresh 
Type: Update 
Category: LSC 
Subject: LSC model updates 

The LSC model has been updated:

Binary outputs to control whitening filter switching

We now take the filter state bit from the first filter bank in all RF PD I/Q filter banks (AS55_I, REFL11_Q, etc) as the controls for the binary analog whitening switching on the RF PD I/Q inputs. The RF_PD part was also modified to output this control bit. The bits from the individual PDs are then combined into the various words that are written to the Contec BO part.

Channel mapping updated/fixed to reflect wiring specification

Yesterday Suresh posted an updated LSC wiring diagram, with correct channel assignments for the RF PD I/Q and DC inputs.  Upon inspection of the physical hardware we found that some of LSC the wiring was incorrect, with I/Q channels swapped, and some of the PDs in the wrong channels.  We went through and fixed the physical wiring to reflect the diagram.  This almost certainly will affect the EPICS settings for some of the input channels, such as offsets and RD rotations.  We should therefore go through all of the RF inputs and make sure everything is kosher.

I also fixed all of the wiring in the LSC model to also reflect the diagram.

Once this was all done, I rebuilt and restarted the LSC model, and confirmed that the anti-whitening filter banks in the PD input filter modules were indeed switching the correct bits.  I'll next put together a script to confirm that the LSC PD whitening is switching as it should.

 

ELOG V3.1.3-